# PriVi: Towards A General-Purpose Video Model For Primate Behavior In The Wild

### v1

@@ 5.1 PriVi improves performance over human-centric and in-domain pretraining
VideoPrism-g [61] $31.5$
Ours
PriVi 86.74 62.75 39.25 58.16 75.43 $33.99$ 95.58 $71.30$
PriVi+CID 85.01 62.96 40.00 59.29 76.42 38.57 96.02 75.14
@@
To measure the effect of our YouTube relevance filtering, we also compare to the baseline of randomly selected YouTube videos YT-Random. Our relevance filtering considerably improves performance (Table˜2, YT-Filtered vs. YT-Random). To our surprise, even no relevance filtering still mostly outperforms the V-JEPA baseline, suggesting that even noisy, partly primate-centric data is an improvement over VideoMix2M.
Broad PriVi pretraining outperforms in-domain continual pretraining
If pretraining on in-domain data is feasible, is it still beneficial to use PriVi or should one simply train on the in-domain data? We pretrain on videos from ChimpACT and PanAf500 without labels and compare with PriVi. Even though in-domain pretraining works, it is consistently outperformed by broad PriVi pretraining (Table˜2). We also do not observe transfer learning between ChimpACT and PanAf500, while PriVi performs well on both datasets.
In-domain unlabeled video data boosts pretraining performance further
PriVi includes only broad, domain-related videos of primates. To measure how much performance increase one could expect from including unlabeled in-domain data in the pretraining, we perform continual pretraining on the target dataset (ChimpACT, PanAf500, BaboonLand or ChimpBehave) after pretraining on PriVi. Note that this pre-training is still entirely unsupervised. We use only detected, not ground-truth bounding boxes. Continual pertraining improves performance further, especially on class-balanced metrics, with $2.68\,\%$ mAP on ChimpACT and $7.34\,\%$ balanced accuracy on PanAf500 (Table˜3).
@@ 5.2 PriVi outperforms prior methods
Next, we compare our method against current state-of-the-art methods. Pretraining on PriVi followed by in-domain continual pretraining (PRIMA+CID) surpasses the state of the art on all four datasets across both class-balanced and unbalanced metrics (Section˜5.1). Our lightweight frozen attentive classifier using 220 k parameters outperforms large specialist models like ChimpVLM with 167 M parameters and full finetuning of human-centric models like VideoMAEv2.
@@
When comparing V-JEPA pretrained on only human-centric data with pretraining on PriVi only, PriVi brings considerable performance improvements across all four datasets.
@@ 5.3 Ablations
Finally, we establish necessity of each component by ablating various design choices for pretraining and attentive classifier. Attentive classifiers without aggressive downprojection have orders of magnitude more parameters than our approach. To ensure a fair comparison, we explore learning rates between $1\text{\times}{10}^{-3}$ and $5\text{\times}{10}^{-5}$ and perform early stopping for each ablation.
@@
During pretraining, we crop videos to detected primate bounding boxes, because we assume that it is inefficient to spend computational resources on learning representations for background patches. Indeed, primate-centric cropping improves performance considerably on every metric compared with training on full video frames (Table˜5; w/o PrimateCrop).
@@
Removing the linear downprojection (w/o downproject), reducing from three to one layers (single layer), and using one instead of $C$ class tokens (single cls token) reduces performance across all metrics (Table˜5), even though the effect of $C$ class tokens is small. Even using only one layer still achieves impressive performance with only 120 k parameters, highlighting the quality of representation produced of our pretrained model.
@@
A single cross-attention layer (Cross-Att.) is commonly used for frozen evaluation to demonstrate representation quality [4, 51]. While this setup performs worse than our improved attentive classifier with more layers (except on PanAf500 balanced accuracy), it still shows good performance, demonstrating good feature quality.
@@
Table 5: All components of our architecture contribute positively. We ablate (a) pretraining and (b) frozen classification design decision and (c) compare our classifier against other common attentive classifiers. All models with pretraining on PriVi only, results reported on validation sets.
Head Params ChimpACT PanAf500
@@ 6 Conclusion
We introduced both PriVi, a large-scale pretraining dataset for primate behavior recognition, along with a framework for effectively utilizing unlabeled data for primate behavior recognition. We assembled PriVi from behavioral research footage spanning various species and settings, as well as YouTube videos of primates for diversity. We propose a simple, scalable data curation pipeline based on CLIP embeddings and zero-shot detections. We pretrain a V-JEPA video model using our data and use it to verify that both data sources as well as the curation pipeline contribute positively for primate behavior recognition, surpassing the state-of-the-art on all four primate behavior datasets. Thus, domain-related pretraining yields considerable performance over using a foundation model pretrained on mostly human-centric data directly.
Our proposed animal behavior recognition framework consists of (a) self-supervised pretraining of a video model on large-scale diverse animal-centric videos, (b) optional in-domain continual pretraining on the specific target domain and (c) frozen evaluation using a narrow, but deep attentive classifier. This framework has several desirable properties: Domain-related pretraining yields good behavior recognition performance across a wide variety of datasets; continual in-domain pretraining further boosts performance for individual datasets without requiring labeled data; and our frozen evaluation setup performs well even on few labeled samples. We hope that our framework and dataset will serve as a strong starting point for the community to move from specialist models for each dataset towards unified primate behavior models, which could open new avenues for research in ethology, ecology, and conservation biology.
@@
Even though our datasets sets a new baseline for diversity in primate behavior datasets, it still captures only 11 different research setups. Due to the limited availability of labeled datasets, we could only evaluate on chimpanzees and baboons and all of our experiments currently utilize ground truth annotated bounding boxes.
Future work
A relevant avenue for future work would be transferring our framework to other orders and suborders of mammals and eventually unifying models across mammals, as soon as sufficient pretraining data is available. Combining our approach with language models could prove beneficial for very small datasets, as it would allow in-context learning and utilization of ethograms. Further research into efficient models for real-world deployment of primate behavior recognition is needed, e. g. by distilling large video models using datasets like PriVi.

### v2

@@ 4.2 Architecture and Pretraining Details
Our Classifier
(V-JEPA) $36.33$ $55.50$ $82.96$ $56.69$ $74.91$ $26.99$ $94.99$ $68.41$
PriVi 39.25 58.16 86.74 62.75 75.43 33.99 95.58 $71.30$
PriVi + CID 40.00 59.29 85.01 62.96 76.42 38.57 96.02 75.14
Figure 5: Our approach scales favorably with fewer labels. Error bars are 95 % confidence intervals estimated over three different subsets of the dataset. Results on validation sets.
Continual Pretraining
Following V-JEPA [4], we choose a ViT-L model with video input for $E$ and a narrow 12-layer transformer for the V-JEPA predictor $P$. $E$ has 304 M parameters, $P$ has 22 M parameters. We initialize the weights from a checkpoint pretrained on VideoMix2M, consisting of Kinetics, SomethingSomethingv2, and HowTo100M, and perform pretraining for 75 k steps with a batch size of 80. This corresponds to eight epochs on the PriVi dataset. As the initialized weights have already been cosine annealed, we only perform warmup and train with a constant small learning rate of $1.5\text{\times}{10}^{-5}$, following Singh et al. [52]. Training takes 11 h on a single node with 4 A100 GPUs.
Continual In-Domain Pretraining (CID)
PriVi includes only broad, domain-related videos of primates. To evaluate how much performance increase one could expect from unlabeled in-domain data, we continue pretraining on the target dataset (ChimpACT, PanAf500, BaboonLand or ChimpBehave) after pretraining on PriVi. Note that this pre-training is still entirely unsupervised, we do not use ground-truth bounding boxes. Chunking and primate detection is performed as for R&O (Sec.˜3.2) and we exclude the test set from pretraining except for ChimpBehave which does not have a dedicated test set. We pretrain as we do on PriVi but only train for 10k steps.
Attentive Classifier
On ChimpACT and PanAf500, we report performance at the best validation accuracy, for ChimpBehave and BabboonLand, we report performance at the end of training. To reduce noise from small datasets, we calculate class scores on three views per test sample and train five attentive classifiers and report the average. We always sample 16 frames uniformly from the input. We use cross-entropy loss except for BaboonLand where standard protocol is EQL loss [54].
5 Results
We first explore how varying pretraining data impacts behavior recognition for a fixed model (Sec.˜3.3) and classifier (Sec.˜3.4) architecture. We initialize using V-JEPA weights pretrained on VideoMix2M and then pretrain with only YT-Filtered, only R&O and both combined (i. e. PriVi). Both YT-Filtered and R&O individually improve performance over standard V-JEPA on both PanAf and ChimpACT, even though R&O shows smaller gains (Tab.˜2). R&O contains camera trap recordings of chimpanzees, but no recordings of zoo-housed chimpanzees, which is mirrored in the larger performance improvements on PanAf500 compared to ChimpACT (Tab.˜2, V-JEPA vs. R&O). Combining YT-Filtered and R&O to PriVi yields further gains on all metrics (Tab.˜2), underscoring the importance of all parts of the dataset. Figure˜4 shows example predictions of our model.
@@ 5.1 PriVi improves performance over human-centric and in-domain pretraining
To measure the effect of our YouTube relevance filtering, we also compare to the baseline of randomly selected YouTube videos YT-Random. Our relevance filtering considerably improves performance (Tab.˜2, YT-Filtered vs. YT-Random). To our surprise, even no relevance filtering still mostly outperforms the V-JEPA baseline, suggesting that even noisy, partly primate-centric data is an improvement over VideoMix2M.
Broad PriVi pretraining outperforms in-domain pretraining
We continue V-JEPA pertraining on only videos from ChimpACT and PanAf500 (without labels and test samples) and compare with PriVi. Even though this in-domain pretraining works, it is consistently outperformed by broad PriVi pretraining (Tab.˜2). We do not observe transfer learning between ChimpACT and PanAf500 (Tab.˜2, gray values), while PriVi performs well on both.
Continual in-domain pretraining (CID) boosts performance further
Pretraining first on PriVi and then on the target datasets consistently improves performance, except for accuracy on PanAf500, where we see mixed results, likely due to the availability of chimpanzee camera trap recordings in PriVi (Tab.˜2 & 4.2; PriVi vs. PriVi + CID).
@@ 5.2 PriVi outperforms prior methods
We compare our method against current state-of-the-art methods. Pretraining on PriVi followed by in-domain continual pretraining (PriVi + CID) surpasses the state of the art on all four datasets across both class-balanced and unbalanced metrics (Sec.˜4.2). Our lightweight frozen attentive classifier using 220 k parameters outperforms large specialist models like ChimpVLM with 167 M parameters and full finetuning of human-centric models like VideoMAEv2.
@@
When comparing V-JEPA pretrained on only human-centric data with pretraining on PriVi, PriVi brings considerable performance improvements across all datasets (Sec.˜4.2, Our Classifier).
5.3 Labeled Data Efficiency
Next, we evaluate how well our method works on smaller datasets. We produce subsets of ChimpACT and PanAf500 with only 50 %, 25 %, and 10 % of training data, always including or excluding full sequences. PanAf500 contains 400 training video sequences, while ChimpACT contains only 127, making the task more challenging on ChimpACT. We compare against X3D [17] as baseline.
We find that on PanAf500, our method scales very favorably with fewer labels, losing only 4.4 % accuracy from 100 % to 10 % training data. Ours with 10 % labeled data outperforms X3D with 100 % labeled data. X3D also scales very well up to 25 % without losing accuracy, but drops in performance at 10 %. On ChimpACT, both methods lose more performance when reducing labeled data, however our method at 25 % still outperforms X3D at 100 %. See Fig.˜5.
@@ 5.4 Ablations
Finally, we establish necessity of each component by ablating various design choices for pretraining and attentive classifier. Attentive classifiers without downprojection have far more parameters than our approach. To ensure a fair comparison, we explore learning rates between $1\text{\times}{10}^{-3}$ and $5\text{\times}{10}^{-5}$ and perform early stopping for each ablation.
@@
During pretraining, we crop videos to detected primate bounding boxes, because we assume that it is inefficient to spend computational resources on learning representations for background patches. Indeed, primate-centric cropping improves performance considerably on every metric compared with training on full video frames (Tab.˜4, w/o primate crop).
@@
Removing the linear downprojection, reducing from three to one layers, and using one instead of $C$ class tokens reduces performance across all metrics (Tab.˜4), even though the effect of $C$ class tokens is small. Even using only one layer still achieves impressive performance with only 120 k parameters, highlighting the quality of representation produced by our pretrained model.
@@
A single cross-attention layer (Cross-Att.) is commonly used for frozen evaluation to demonstrate representation quality [4, 63, 53]. While this setup performs worse than our improved attentive classifier with more layers (except on PanAf500 balanced accuracy), it still shows good performance, demonstrating high feature quality.
@@
Table 4: All components of our architecture contribute positively. We ablate (a) pretraining and (b) classifier design decision and (c) compare our classifier against other common attentive classifiers. All models with pretraining on PriVi only, results reported on validation sets.
Train. Params ChimpACT PanAf500
@@ 6 Conclusion
We introduced PriVi, a large-scale primate-centric pretraining dataset, and a simple framework for leveraging unlabeled video to improve primate behavior recognition. PriVi combines curated research footage from diverse species and settings with filtered YouTube videos, assembled through a scalable pipeline using CLIP-based relevance scoring and zero-shot primate detection. Pretraining V-JEPA on this dataset consistently boosts performance and surpasses prior work across all four benchmark datasets, showing the clear advantage of domain-related pretraining over human-centric alternatives.
Our proposed animal behavior recognition framework consists of (a) self-supervised pretraining on large-scale diverse animal-centric videos, (b) optional in-domain continual pretraining (CID) on the specific target domain and (c) frozen evaluation using a narrow, but deep attentive classifier. This framework has several desirable properties: Domain-related pretraining yields good behavior recognition performance across a wide variety of datasets; CID further boosts performance for individual datasets without requiring labeled data; and our frozen evaluation performs well even on few labeled samples. We hope that our framework and dataset will serve as a strong starting point for the community to move from specialist models towards unified primate behavior models, which could open new avenues for research in ethology, ecology, and conservation biology.
@@
Even though our dataset sets a new baseline for diversity in primate behavior datasets, it still captures only 11 different research setups. Due to the limited availability of labeled datasets, we could only evaluate on chimpanzees and baboons and all of our experiments currently utilize ground truth annotated bounding boxes.